# Endomorphism algebras of abelian varieties with large cyclic 2-torsion field over a given field

### v3

@@ 2. Proof of the main theorems
It is a routine calculation to show $\mathcal{O}$ is an order of discriminant $D$ (for an example, see [Vig80, pg. 85 - 86]).
As $m$ is odd we have $\mathbb{Z}[1,mi,j,k]\otimes\mathbb{Z}_{2}=\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{2}$. Thus we may assume $\mathcal{O}^{\prime}$ contains $\mathbb{Z}[1,i,j,k]$. The discriminant of $\mathbb{Z}[1,i,j,k]$ equals $4D$ and thus any element contained in $\mathcal{O}^{\prime}$ but not in $\mathbb{Z}[1,i,j,k]$ is of the form $\frac{1}{4}\alpha$ with $\alpha\in\mathbb{Z}[1,i,j,k]$.
@@
Hence it suffices to check elements of the form $\frac{1}{2}(a+bi+cj+dk)$ with $a,b,c,d\in\{0,1\}$. We have $\mathrm{n}(a+bi+cj+dk)\equiv a-2b+c+2d\equiv 0\mod{4}$, from which we deduce $\frac{1}{2}(1+j+k)$, $\frac{1}{2}(1+i+j)$, $\frac{1}{2}(i+k)$ are the only integral such. In order for the discriminant of $\mathcal{O}^{\prime}$ to equal $D$ we see that all of these elements must belong to $\mathcal{O}^{\prime}$. ∎
@@
Moreover, any order $\mathcal{O}^{\prime}$ of even discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$ satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$. Any order $\mathcal{O}^{\prime}$ of discriminant $D\equiv t\mod{4}$ with $t\in\{1,3\}$, which contains $\mathbb{Z}[1,mi,j,k]$, satisfies either $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$ or $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}_{t}\otimes\mathbb{Z}_{2}$.
@@
As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Hence Lemma 3.1 allows us to take $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j+k)\mathbb{Z}+\frac{1}{2}(1+j-k)\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
Let $X=\frac{1}{2}(1+j+k)$, $Y=\frac{1}{2}(1+j-k)$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=1-Y$ and $iZi^{-1}=Z+Y-X$. For $j$ we have $jXj^{-1}=Y$, $jYj^{-1}=X$, $jZj^{-1}=-Z$. Looking at the coefficients, we see the action remains faithful on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. By proceeding as in the proof of Theorem 2.1 we are now done. ∎
@@
As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Lemma 3.2 allows us to take $\operatorname{{\rm End}}(A)$ equal to one of the three given orders. As these orders differ by permuting $i,j$ and $k$, we only give details for the case $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j)\mathbb{Z}+k\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
Let $X=\frac{1}{2}(1+j)$, $Y=k$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=-Y$ and $iZi^{-1}=Z-Y$. For $j$ we have $jXj^{-1}=X$, $jYj^{-1}=-Y$, $jZj^{-1}=-Z$. Looking at the coefficients, we see $j$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ whereas $i$ and $k$ act by the same involution. Proceeding as in the proof of Theorem 2.1, we find at least one of $\mathbb{Q}(i)\cong\mathbb{Q}(\sqrt{D/m})$ and $\mathbb{Q}(k)\cong\mathbb{Q}(\sqrt{-D})$ is contained in $\operatorname{{\rm End}}_{F}^{0}(A)$. ∎

### v4

@@ 2. Proof of the main theorems
It is a routine calculation to show $\mathcal{O}$ is an order of discriminant $D$ (for an example, see [Vig80, p. 85 - 86]).
As $m$ is odd we have $\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{(2)}=\mathbb{Z}[1,mi,j,k]\otimes\mathbb{Z}_{(2)}\subseteq\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}$. There exists a lattice $\mathcal{O}^{\prime\prime}$ such that $\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}$ and $\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(p)}=\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{(p)}$ for every odd prime $p$ [Voi21, Thm. 9.4.9]. Moreover, $\mathcal{O}^{\prime\prime}$ is in fact an order since every localisation is [Voi21, Lem. 10.2.10]. As $\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{(p)}$ is contained in $\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(p)}$ for every prime $p$, we deduce $\mathbb{Z}[1,i,j,k]$ is contained in $\mathcal{O}^{\prime\prime}$ [Voi21, Cor. 9.4.7].
The discriminant of $\mathbb{Z}[1,i,j,k]$ equals $4D$. As the discriminant of an order is determined locally [Voi21, 15.2.13], we deduce the discriminant of $\mathcal{O}^{\prime\prime}$ equals $D$. Applying [Voi21, Lem. 15.2.5], we see that any element contained in $\mathcal{O}^{\prime\prime}$ but not in $\mathbb{Z}[1,i,j,k]$ is of the form $\frac{1}{4}\alpha$ with $\alpha\in\mathbb{Z}[1,i,j,k]$.
@@
Hence it suffices to check elements of the form $\frac{1}{2}(a+bi+cj+dk)$ with $a,b,c,d\in\{0,1\}$. We have $\mathrm{n}(a+bi+cj+dk)\equiv a-2b+c+2d\equiv 0\mod{4}$, from which we deduce $\frac{1}{2}(1+j+k)$, $\frac{1}{2}(1+i+j)$, $\frac{1}{2}(i+k)$ are the only integral such. In order for the discriminant of $\mathcal{O}^{\prime\prime}$ to equal $D$ we see that all of these elements must belong to $\mathcal{O}^{\prime\prime}$. Whence $\mathcal{O}\otimes\mathbb{Z}_{(2)}=\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}$. ∎
@@
Moreover, any order $\mathcal{O}^{\prime}$ of even discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$ satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}\otimes\mathbb{Z}_{(2)}$. Any order $\mathcal{O}^{\prime}$ of discriminant $D\equiv t\mod{4}$ with $t\in\{1,3\}$, which contains $\mathbb{Z}[1,mi,j,k]$, satisfies either $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}\otimes\mathbb{Z}_{(2)}$ or $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}_{t}\otimes\mathbb{Z}_{(2)}$.
@@
If $L=K$, then there is nothing left to show. Thus let us suppose $L\neq K$. As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Thm. 3.4] apply to show $\operatorname{{\rm Gal}}(L/K)\leq C_{2}^{2}$ and a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we shall consider a $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ reduced modulo $2$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Hence Lemma 3.1 allows us to take $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}=\mathbb{Z}_{(2)}+\frac{1}{2}(1+j+k)\mathbb{Z}_{(2)}+\frac{1}{2}(1+j-k)\mathbb{Z}_{(2)}+\frac{1}{2}(i+k)\mathbb{Z}_{(2)}$.
Let $X=\frac{1}{2}(1+j+k)$, $Y=\frac{1}{2}(1+j-k)$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=1-Y$ and $iZi^{-1}=Z+Y-X$. For $j$ we have $jXj^{-1}=Y$, $jYj^{-1}=X$, $jZj^{-1}=-Z$. Looking at the coefficients, we see the action remains faithful on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$.
Recall that the natural map $\operatorname{{\rm End}}(A)\rightarrow\operatorname{{\rm End}}(A[2])$ induces an injection $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}(A[2])$. The group $\operatorname{{\rm Gal}}(\overline{K}/K(A[2]))$ acts trivially on $\operatorname{{\rm End}}(A[2])$ and hence trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ as well. Since $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, we find $L\subseteq K(A[2])$. ∎
@@
If $L=K$, then there is nothing left to show. Thus let us suppose $L\neq K$. As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Thm. 3.4] apply to show $\operatorname{{\rm Gal}}(L/K)\leq C_{2}^{2}$ and a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we shall consider a $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ reduced modulo $2$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Lemma 3.2 allows us to take $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ equal to one of the three given orders tensored by $\mathbb{Z}_{(2)}$. As these orders differ by permuting $i,j$ and $k$, we only give details for the case $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}=\mathbb{Z}_{(2)}+\frac{1}{2}(1+j)\mathbb{Z}_{(2)}+k\mathbb{Z}_{(2)}+\frac{1}{2}(i+k)\mathbb{Z}_{(2)}$.
Let $X=\frac{1}{2}(1+j)$, $Y=k$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=-Y$ and $iZi^{-1}=Z-Y$. For $j$ we have $jXj^{-1}=X$, $jYj^{-1}=-Y$, $jZj^{-1}=-Z$. Looking at the coefficients, we see $j$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ whereas $i$ and $k$ act by the same involution.
We thus find that $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ through a quotient group $\operatorname{{\rm Gal}}(L^{\prime}/K)$ of order at most 2. As $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}(A[2])$, we find $L^{\prime}\subseteq K(A[2])$. The above calculation shows the action of $\operatorname{{\rm Gal}}(L^{\prime}/K)$ is via conjugation by one of either $i$ or $k$. Hence, we find at least one of $\mathbb{Q}(i)\cong\mathbb{Q}(\sqrt{D/m})$ and $\mathbb{Q}(k)\cong\mathbb{Q}(\sqrt{-D})$ is contained in $\operatorname{{\rm End}}_{F}^{0}(A)$, see the proof of [DR04, Thm. 3.4]. ∎